# Rough Sets for Explainability of Spectral Graph Clustering

### v1

@@ 3.1 Review of literature on handling boundary data in clustering via rough sets
One of the methods used in the past to determine a set of objects that unambiguously belong to a cluster and another larger set that possibly contains it was developed based on the so-called credal partitioning, or evidential clustering, [Masson:2008:ECM]. Interestingly, this approach is suitable for adding partial a priori knowledge about cluster membership [Antoine:2012:CECM].
@@
Another method was proposed by [Whang:2015:nonexhaus]. It is based on an extension of weighted $k$-means to make it both overlapping (so that there are boundary objects) and non-exhaustive (to take care of outliers). Parameters control the maximum overlapping level and the share of outliers. It turns out to be applicable to graph cut approximations.
@@
While there seems to be no direct study on the explainability of rough set-based clustering, explainability itself was a part of various rough set-based prediction methodologies like that of [Cao:2022]. [GrzegorowskiJSMS:2023] used the clustering and rough set methods to obtain human-readable cluster descriptions, emphasizing each cluster’s most discernible characteristics. [Singh2017] propose to use reducts (an important notion used in rough set theory) to explain the existence of clusters and to clarify why two clusters are different.
@@
[WANG:201859] apply the graph spectral clustering to non-textual data with the intent of semantic cluster content description based on fuzzy methodology. The authors concentrate on appropriate feature selection to get the best possible cluster separation. [CEKIK2020113691] apply rough set theory to find the best features in short textual documents in the context of classification.
@@
We already know that replacing the Euclidean distance with kernel-induced metric makes it possible to cluster the objects that belong to non-convex clusters. This gave birth to rough-kernel clustering algorithms. Let us mention a few of them: [Zhu-2008], [Das-2009], [hu2010kernelized], and [Tripathy-2012].
@@ 3.2 Our approach to handling boundary data in clustering inspired by rough sets
The difference of our approach to the problems with distortions of clusterings lies essentially in insisting that the clustering results need to be explainable in the textual domain. As we already wrote in [Plosone2025], there is a big difference between describing a cluster with a vector of most frequent words occurring therein and the explanation of the cluster. The explanation means that the cluster membership of documents in a cluster is justified by its cosine similarity to cluster prototype (centroid) in the term-vector space. And these two issues are not equivalent. In [Plosone2025] we demonstrated that the traditional $L$-based and $N$-based GSC methods can be translated via our proprietary $K$-embedding and $\mathcal{M}$-embedding can be approximately equivalently translated to term-vector space representation and thus be explainable textually by term vectors. None of the aforementioned publications seems to care about it. Only cluster descriptions are considered.
What is more, even if some methodologies are based on GSC, like [Yue:2023], and their intermediate results rely on cosine similarity, their approach does not support the explainability, as they rely on GSC clustering co-occurrences in clusterings from various clustering methods rather than on derivation of clusters that are explainable, which makes the connection to the original document similarity is lost.
@@ 3.3 Theoretical Background
Let us consider a bit theoretical clustering of objects that have nonzero similarities within clusters and zero similarities between clusters. Let the similarity matrix be sorted in such a way that it looks like
@@
whereby the symmetric matrices $S_{[n_{i-1}+1:n_{i}]}$, $i=1,\dots,k$ are the similarity matrices within the clusters $C_{1},\dots,C_{k}$ (with zero diagonals), $0<n_{1}<n_{2}<\dots<n_{k}$, $n_{k}$ being the number of objects.
@@
In order to perform GSC clustering, one takes $k$ eigenvectors of $L(S)$ assigned to the lowest eigenvalues, whereby any eigenvalue of $L(S)$ is provably non-negative.
Our $L(S)$ has $k$ eigenvalues equal to zero with corresponding eigenvector $\mathbf{v}_{i}=(v_{i,1},\dots,v_{i,n_{k}})^{T}$, $i\in 1:k$ such $v_{i,\ell}=1/\sqrt{n_{i}-n_{i-1}}$ for $\ell\in n_{i-1}+1:n_{i}]$ and $v_{i,\ell}=0$ for $\ell\not\in n_{i-1}+1:n_{i}]$. Clearly $\|\mathbf{v}_{i}\|=1$, as required by GSC methodology.
@@
Now $L(S^{\prime})$ has $k+1$ eigenvalues equal to 0 and same corresponding eigenvalues as $L(S)$ plus the eigenvector $\mathbf{v}_{k+1}=(0,0,\dots,0,1)^{T}$. Now we have the problem. GSC can pick up the right eigenvectors (the first $k$) and we are lucky, or the k+1k+1st eigenvector will be taken for clustering and then we get a problem. The last object may have coordinates close enough to an intrisic cluster center and then we are lucky, or it may be too far away so that the clustering process singles it out as one separate cluster while collapsing two intrinsic clusters to a single one.
This may happen as follows: In case we take the eigenvector $\mathbf{v}_{k+1}$ for embedding the objects, then one of the intrisic clusters will have the center at $\mathbf{0}$ (coordinate origin), and the other at $(0,...,1/\sqrt{n_{i}-n_{i-1}},....0)^{T}$. So he distance of the dissimilar object to the intrinsic cluster centers will amount either to $1$ or to $\sqrt{1+\frac{1}{n_{i}-n_{i-1}}}$ whereas the distance between intrinsic clusters will amount to either $1/\sqrt{n_{i}-n_{i-1}}$ or to $\sqrt{\frac{1}{n_{i^{\prime}}-n_{i^{\prime}-1}}+\frac{1}{n_{i}-n_{i-1}}}$. With increasing cluster sizes the distance between trueclusters will decrease down to zero, while the dissimilar element’s distance will get closer to 1. At a point it will be more advantageous for $k$-means to separate the dissimilar element as a single cluster and to collapse two intrinsic clusters. With increasing number of elements dissimilar to all the other elements, the situation will get worse.
@@
In real settings, this situation of a pure dissimilar object may happen rarely, but instead low similarity objects may exist so that the behaviour of the clustering process will approximate the mentioned one.
@@ 4.4 Cluster Explanations
Though this paper does not deal with the textual explanation per se (an interested Reader shall refer to [Plosone2025]), nonetheless it is worthy to have a look at changes induced by the spurious data filtering technique applied. Note that the keywords presented below are not the most frequent ones, but rather are selected by an explanation process of cluster centroids. We observe, for example, that after the application of the technique, the $L$-based cluster explanations appear to be more reasonable.
@@ 5 Conclusions
The broad set of methods of Graph Spectral Clustering, until recently, has been considered as black-box clustering methods, without any possibility of explanation of clustering results. Only recently, some methods of explanation have been proposed, for example in [Plosone2025]. However, under practical settings, the GSC methods produced high discrepancy between the results. In this paper, we proposed a new approach to explainability, or rather an amendment to the just mentioned proposals. The amendment consists in removing a part of data with low discernability of cluster membership versus non-membership and concentrate only on the high discernability elements. The explanation to these clusters will be the explainability foundation. In this way, various clustering methods get closer to one another in their results so that finally the intersection between them can be considered a final explanation.
@@
Last but not least the issue of algorithm convergence was studied. Frequently, noise in the data is addressed by seeking intersection of results of various clustering methods. A simplistic application of such an approach may be disastrous as various algorithms have different built-in clustering target functions. Only those algorithms can be applied in this context that are able to detect same clustering when the data is ideally separated, that is without boundary elements. We have demonstrated that a number of GSC algorithms have this convergence property, and separate publications ([Plosone2025]) demonstrate proofs of equivalence or near equivalence.

### v2

@@ 3.1 Review of literature on handling boundary data in clustering via rough sets
One of the methods used in the past to determine a set of objects that unambiguously belong to a cluster and another larger set that possibly contains it was developed based on the so-called credal partitioning, or evidential clustering, [26]. Interestingly, this approach is suitable for adding partial a priori knowledge about cluster membership [1].
@@
Another method was proposed by [53]. It is based on an extension of weighted $k$-means to make it both overlapping (so that there are boundary objects) and non-exhaustive (to take care of outliers). Parameters control the maximum overlapping level and the share of outliers. It turns out to be applicable to graph cut approximations.
@@
While there seems to be no direct study on the explainability of rough set-based clustering, explainability itself was a part of various rough set-based prediction methodologies like that of [5]. [14] used the clustering and rough set methods to obtain human-readable cluster descriptions, emphasizing each cluster’s most discernible characteristics. [44] propose to use reducts (an important notion used in rough set theory) to explain the existence of clusters and to clarify why two clusters are different.
@@
[52] apply the graph spectral clustering to non-textual data with the intent of semantic cluster content description based on fuzzy methodology. The authors concentrate on appropriate feature selection to get the best possible cluster separation. [6] apply rough set theory to find the best features in short textual documents in the context of classification.
@@
We already know that replacing the Euclidean distance with kernel-induced metric makes it possible to cluster the objects that belong to non-convex clusters. This gave birth to rough-kernel clustering algorithms. Let us mention a few of them: [59], [8], [17], and [48].
@@ 3.2 Our approach to handling boundary data in clustering inspired by rough sets
The difference of our approach to the problems with distortions of clusterings lies essentially in insisting that the clustering results need to be explainable in the textual domain. As we already wrote in [45], there is a big difference between describing a cluster with a vector of most frequent words occurring therein and the explanation of the cluster. The explanation means that the cluster membership of documents in a cluster is justified by its cosine similarity to cluster prototype (centroid) in the term-vector space. And these two issues are not equivalent. In [45] we demonstrated that the traditional $L$-based and $N$-based GSC methods can be translated via our proprietary $K$-embedding and $\mathcal{M}$-embedding can be approximately equivalently translated to term-vector space representation and thus be explainable textually by term vectors. None of the aforementioned publications seems to care about it. Only cluster descriptions are considered.
What is more, even if some methodologies are based on GSC, like [56], and their intermediate results rely on cosine similarity, their approach does not support the explainability, as they rely on GSC clustering co-occurrences in clusterings from various clustering methods rather than on derivation of clusters that are explainable, which makes the connection to the original document similarity is lost.
@@ 3.3 Theoretical Background
Let us consider an idealized case in which objects have nonzero similarities within clusters and zero similarities between clusters. Suppose also that the objects are arranged such that the similarity matrix has the block-diagonal structure, i.e.
@@
whereby the symmetric matrices $S_{[N_{i-1}+1:N_{i}]}$, $i=1,\dots,k$ are the similarity matrices within the clusters $C_{1},\dots,C_{k}$ (with zero diagonals), $N_{0}=0.N_{i}=N_{i-1}+n_{i}$ for $i\geq 1$, $0<n_{1}<n_{2}<\dots<n_{k}$, $n_{k}$ being the number of objects.
@@
To perform GSC clustering, one takes $k$ eigenvectors of $L(S)$ assigned to the lowest eigenvalues, whereby any eigenvalue of $L(S)$ is provably non-negative as $L(S)$ is a Laplacian. Let us denote with $\mathbf{v^{(i)}}$ a vector with all elements equal to $\frac{1}{\sqrt{n_{i}}}$. It is of unit length and it is an eigenvector of $L(S_{i})$ which corresponds to the lowest eigenvalue of it (that is 0). What is more, as observed in [29], due to block-diagonality, our $L(S)$ has $k$ eigenvalues equal to zero with corresponding mutually orthogonal eigenvectors $\mathbf{v}_{i}=(v_{i,1},\dots,v_{i,n_{k}})^{T}$, $i\in 1:k$ such that $v_{i,\ell}={v^{(i)}}_{\ell-N_{i-1}}$ for $\ell\in[N_{i-1}+1:N_{i}]$ and $v_{i,\ell}=0$ for $\ell\not\in[N_{i-1}+1:N_{i}]$.
@@
Now $L(S^{\prime})$ has $k+1$ eigenvalues equal to 0 and same corresponding eigenvalues as $L(S)$ plus the eigenvector $\mathbf{v}_{k+1}=(0,0,\dots,0,1)^{T}$. We have the problem. GSC can pick up the right eigenvectors (the first $k$) and we are lucky, or the k+1k+1st eigenvector will be taken for clustering and then we get a problem. The last object may have coordinates close enough to an intrinsic cluster center and then we are lucky, or it may be too far away so that the clustering process singles it out as one separate cluster while collapsing two intrinsic clusters to a single one.
This may happen as follows: In case we take the eigenvector $\mathbf{v}_{k+1}$ for embedding the objects, then one of the intrinsic clusters will have the center at $\mathbf{0}$ (coordinate origin), and the other at $(0,...,1/\sqrt{n_{i}},....0)^{T}$. So the distance of the dissimilar object to the intrinsic cluster centers will amount either to $1$ or to $\sqrt{1+\frac{1}{n_{i}}}$ whereas the distance between intrinsic clusters will amount to either $1/\sqrt{n_{i}}$ or to $\sqrt{\frac{1}{n_{i^{\prime}}}+\frac{1}{n_{i}}}$. With increasing cluster sizes the distance between true clusters will decrease down to zero, while the dissimilar element’s distance will get closer to 1. At a point it will be more advantageous for $k$-means to separate the dissimilar element as a single cluster and to collapse two intrinsic clusters. With increasing number of elements dissimilar to all the other elements, the situation will get worse.
@@
In real settings, this situation of pure block-diagonality and of a pure dissimilar object may happen rarely, but instead low similarity objects may exist so that the behavior of the clustering process will approximate the mentioned one. This is easily seen from Theorem 2 formulated in [29] implying that given slight noise, the eigenvectors and hence the clustering behave in a similar manner.  [^4] A deeper analysis of such noisy conditions may be found in [34] as well as in [24]. Hence our proposed method will work also.
@@ 4.4 Cluster Explanations
Though this paper does not deal with the textual explanation per se (an interested Reader shall refer to [45]), nonetheless it is worthy to have a look at changes induced by the spurious data filtering technique applied. Note that the keywords presented below are not the most frequent ones, but rather are selected by an explanation process of cluster centroids. We observe, for example, that after the application of the technique, the $L$-based cluster explanations appear to be more reasonable.
@@ 5 Conclusions
The broad set of methods of Graph Spectral Clustering, until recently, has been considered as black-box clustering methods, without any possibility of explanation of clustering results. Only recently, some methods of explanation have been proposed, for example in [45]. However, under practical settings, the GSC methods produced high discrepancy between the results. In this paper, we proposed a new approach to explainability, or rather an amendment to the just mentioned proposals. The amendment consists in removing a part of data with low discernability of cluster membership versus non-membership and concentrate only on the high discernability elements. The explanation to these clusters will be the explainability foundation. In this way, various clustering methods get closer to one another in their results so that finally the intersection between them can be considered a final explanation.
@@
Last but not least the issue of algorithm convergence was studied. Frequently, noise in the data is addressed by seeking intersection of results of various clustering methods. A simplistic application of such an approach may be disastrous as various algorithms have different built-in clustering target functions. Only those algorithms can be applied in this context that are able to detect same clustering when the data is ideally separated, that is without boundary elements. We have demonstrated that a number of GSC algorithms have this convergence property, and separate publications ([45]) demonstrate proofs of equivalence or near equivalence.